# AgentFuzzer: Generic Black-Box Fuzzing for Indirect Prompt Injection against LLM Agents

### v2

@@ 5.3 Ablation Study
We perform an ablation study on AgentDojo to isolate the impact of each of the three core components of AgentFuzzer: the initial corpus of adversarial prompt templates, the adaptive seed scoring strategy, and the MCTS-based seed selection. Specifically, we (1) replace our initial corpus with the handcrafted baseline prompts from AgentDojo, (2) substitute the adaptive seed scoring and MCTS-based seed selection to uniform random seed selection. As shown in Figure 3, AgentFuzzer significantly outperforms the ablated versions. Notably, when the initial corpus is replaced by the baseline prompts, the overall success rate plateaus after approximately four iterations, demonstrating both reduced performance and limited potential compared to our curated initial corpus. Furthermore, without adaptive seed scoring or the MCTS-based seed selection, the fuzzing process shows markedly slower improvement, as it fails to identify and prioritize high-potential seeds. These findings underscore the critical role of all three components in driving AgentFuzzer’s continuous enhancement and superior attack success.
@@ 6 Conclusion
We introduce AgentFuzzer, a novel fuzzing framework designed to systematically conduct indirect prompt injection attacks against blackbox agents with various architectures and tasks. By combining high-quality prompt templates, adaptive seed scoring, and a MCTS-based seed selection algorithm, AgentFuzzer overcomes challenges posed by the black-box nature, architectural complexity, and wide-ranging functionalities of real-world agents. Our empirical results demonstrate that AgentFuzzer not only achieves high attack success rates on established benchmarks and real-world agents but also exhibits strong transferability across unseen tasks and underlying LLMs. By automating the generation and optimization of adversarial prompts, AgentFuzzer highlights critical limitations in existing agent defenses, underlining the urgent need for more robust security measures. We believe AgentFuzzer will serve as a useful foundation for advancing both the understanding of agent-based threats and the development of next-generation security solutions in this rapidly evolving domain.
@@ B.1 Evaluation on Extra Models
While our primary focus has been on commercial black-box LLMs such as GPT, Claude, and Gemini, we also evaluate AgentFuzzer on open-source models. These models often lag behind in long-context understanding, advanced tool usage, reasoning, and planning, which are essential in the challenging agent scenarios. We assess models that support tool calling using the AgentDojo benchmark and report their utility scores (i.e., success rates on benign tasks): Llama3.3-70B-Instruct (Meta AI, 2024) (42%), Qwen2.5-72B-Instruct (Yang et al., 2024) (54%), QwQ-32B (Qwen Team, 2025) (74%), and, for comparison, o3-mini (79%). Based on these results, we conduct further experiments using QwQ-32B. Additionally, the o3-mini checkpoint used in the main text corresponds to an experimental version. We therefore also evaluate the latest available checkpoint, o3-mini-2025-01-31. Results indicate that AgentFuzzer achieves a success rate of 72% on the fuzzing set and 74% on the test set, compared to the baseline handcrafted attack, which achieves 50% and 53%, respectively, as shown in Table 4. These findings underscore AgentFuzzer’s effectiveness, even when applied to strong open-source models.
@@ B.2 Comparison with Additional Baselines
To enhance our baseline comparisons, we included two additional prompt injection baselines from OpenPromptInjection (Liu et al., 2024) and InjecAgent (Zhan et al., 2024). Table 4 reports the attack success rates on the Fuzzing and Test sets across two models, QwQ-32B and o3-mini. These results confirm that AgentFuzzer outperforms state-of-the-art baselines, especially in discovering and exploiting indirect prompt injection vulnerabilities.
Table 4: Attack success rate (ASR) comparison on AgentDojo with AgentFuzzer and three baseline attacks. The two ASRs in each cell represent performance on the Fuzzing task set and Test task set, respectively (i.e., Fuzzing / Test).
@@ B.3 Breakdown on Attack Scenarios
The two benchmarks, AgentDojo and VWA-adv, are designed to evaluate performance across diverse scenarios. We perform additional analysis about the detailed results on different scenarios to provide a comprehensive view of AgentFuzzer’s effectiveness.
AgentDojo consists of various agent tasks grouped into four suites – Slack, Workspace, Travel, and Banking. As shown in Table 5, across all these scenarios, AgentFuzzer consistently achieves a higher success rate compared to the baseline attacks in AgentDojo, demonstrating its robustness and adaptability in different operational environments.
On VWA-adv benchmark, we evaluate performance across two types of adversarial goals: illusioning, which makes it appear to the agent that it is in a different state (e.g., different objects, colors), and goal misdirection, which makes the agent pursue a targeted different goal than the original user goal (e.g., leave a comment). Our results in Table 5 indicate that AgentFuzzer outperforms baseline from the benchmark in both attack goals, confirming its capability to exploit diverse indirect prompt injection vulnerabilities and attack goals effectively, even in challenging goal misdirection tasks.
Table 5: Attack success rates across different scenarios (task suites for AgentDojo, attack goals for VWA-adv) achieved by AgentFuzzer and the baseline attack from the benchmark. The two ASRs in each cell represent performance on the Fuzzing task set and Test task set, respectively (i.e., Fuzzing / Test).
Benchmark Model Scenario AgentFuzzer Benchmark Baseline
@@ B.4 Coverage Curve
The coverage of tasks over fuzzing iterations achieved by AgentFuzzer on VWA-adv benchmark is shown in Figure 4.
Figure 4: Coverage over fuzzing iterations achieved by AgentFuzzer on VWA-adv.

### v4

@@ 5.3 Ablation Study
We perform an ablation study on AgentDojo to isolate the impact of each of the three core components of AgentVigil: the initial corpus of adversarial prompt templates, the adaptive seed scoring strategy, and the MCTS-based seed selection. Specifically, we (1) replace our initial corpus with the handcrafted baseline prompts from AgentDojo, (2) substitute the adaptive seed scoring and MCTS-based seed selection to uniform random seed selection. As shown in Figure 3, AgentVigil significantly outperforms the ablated versions. Notably, when the initial corpus is replaced by the baseline prompts, the overall success rate plateaus after approximately four iterations, demonstrating both reduced performance and limited potential compared to our curated initial corpus. Furthermore, without adaptive seed scoring or the MCTS-based seed selection, the fuzzing process shows markedly slower improvement, as it fails to identify and prioritize high-potential seeds. These findings underscore the critical role of all three components in driving AgentVigil’s continuous enhancement and superior attack success.
@@ 6 Conclusion
We introduce AgentVigil, a novel fuzzing framework designed to systematically conduct indirect prompt injection attacks against blackbox agents with various architectures and tasks. By combining high-quality prompt templates, adaptive seed scoring, and a MCTS-based seed selection algorithm, AgentVigil overcomes challenges posed by the black-box nature, architectural complexity, and wide-ranging functionalities of real-world agents. Our empirical results demonstrate that AgentVigil not only achieves high attack success rates on established benchmarks and real-world agents but also exhibits strong transferability across unseen tasks and underlying LLMs. By automating the generation and optimization of adversarial prompts, AgentVigil highlights critical limitations in existing agent defenses, underlining the urgent need for more robust security measures. We believe AgentVigil will serve as a useful foundation for advancing both the understanding of agent-based threats and the development of next-generation security solutions in this rapidly evolving domain.
@@ B.1 Evaluation on Extra Models
While our primary focus has been on commercial black-box LLMs such as GPT, Claude, and Gemini, we also evaluate AgentVigil on open-source models. These models often lag behind in long-context understanding, advanced tool usage, reasoning, and planning, which are essential in the challenging agent scenarios. We assess models that support tool calling using the AgentDojo benchmark and report their utility scores (i.e., success rates on benign tasks): Llama3.3-70B-Instruct (Meta AI, 2024) (42%), Qwen2.5-72B-Instruct (Yang et al., 2024) (54%), QwQ-32B (Qwen Team, 2025) (74%), and, for comparison, o3-mini (79%). Based on these results, we conduct further experiments using QwQ-32B. Additionally, the o3-mini checkpoint used in the main text corresponds to an experimental version. We therefore also evaluate the latest available checkpoint, o3-mini-2025-01-31. Results indicate that AgentVigil achieves a success rate of 72% on the fuzzing set and 74% on the test set, compared to the baseline handcrafted attack, which achieves 50% and 53%, respectively, as shown in Table 4. These findings underscore AgentVigil’s effectiveness, even when applied to strong open-source models.
@@ B.2 Comparison with Additional Baselines
To enhance our baseline comparisons, we included two additional prompt injection baselines from OpenPromptInjection (Liu et al., 2024) and InjecAgent (Zhan et al., 2024). Table 4 reports the attack success rates on the Fuzzing and Test sets across two models, QwQ-32B and o3-mini. These results confirm that AgentVigil outperforms state-of-the-art baselines, especially in discovering and exploiting indirect prompt injection vulnerabilities.
Table 4: Attack success rate (ASR) comparison on AgentDojo with AgentVigil and three baseline attacks. The two ASRs in each cell represent performance on the Fuzzing task set and Test task set, respectively (i.e., Fuzzing / Test).
@@ B.3 Breakdown on Attack Scenarios
The two benchmarks, AgentDojo and VWA-adv, are designed to evaluate performance across diverse scenarios. We perform additional analysis about the detailed results on different scenarios to provide a comprehensive view of AgentVigil’s effectiveness.
AgentDojo consists of various agent tasks grouped into four suites – Slack, Workspace, Travel, and Banking. As shown in Table 5, across all these scenarios, AgentVigil consistently achieves a higher success rate compared to the baseline attacks in AgentDojo, demonstrating its robustness and adaptability in different operational environments.
On VWA-adv benchmark, we evaluate performance across two types of adversarial goals: illusioning, which makes it appear to the agent that it is in a different state (e.g., different objects, colors), and goal misdirection, which makes the agent pursue a targeted different goal than the original user goal (e.g., leave a comment). Our results in Table 5 indicate that AgentVigil outperforms baseline from the benchmark in both attack goals, confirming its capability to exploit diverse indirect prompt injection vulnerabilities and attack goals effectively, even in challenging goal misdirection tasks.
Table 5: Attack success rates across different scenarios (task suites for AgentDojo, attack goals for VWA-adv) achieved by AgentVigil and the baseline attack from the benchmark. The two ASRs in each cell represent performance on the Fuzzing task set and Test task set, respectively (i.e., Fuzzing / Test).
Benchmark Model Scenario AgentVigil Benchmark Baseline
@@ B.4 Coverage Curve
The coverage of tasks over fuzzing iterations achieved by AgentVigil on VWA-adv benchmark is shown in Figure 4.
Figure 4: Coverage over fuzzing iterations achieved by AgentVigil on VWA-adv.